Find the total energy of a moving object of mass 6.3 π‘˜π‘” and momentum


11.5 Γ— 106 π‘˜π‘”π‘šπ‘ 


βˆ’1


.

1
Expert's answer
2021-12-06T09:48:55-0500

Since the object has a relativistic velocity we need to find the relativistic total energy as follows:


"E=\gamma mc^2=\dfrac{mc^2}{\sqrt{1-\dfrac{v^2}{c^2}}}."

We can find the velocity of the object as follows:


"p=mv,""v=\dfrac{p}{m}=\dfrac{11.5\times10^6\ \dfrac{kg\times m}{s}}{6.3\ kg}=18.25\times10^5\ \dfrac{m}{s}."

Then, we have:


"E=\dfrac{6.3\ kg\times(3\times10^8\ \dfrac{m}{s})^2}{\sqrt{1-\dfrac{(18.25\times10^5\ \dfrac{m}{s})^2}{(3\times10^8\ \dfrac{m}{s})^2}}}=5.67\times10^{17}\ J."